# Neural Network-based Universal Formulas for Control

### v1

@@ Abstract
We study the problem of designing a controller that satisfies an arbitrary number of affine inequalities at every point in the state space. This is motivated by the use of guardrails in autonomous systems. Indeed, a variety of key control objectives, such as stability, safety, and input saturation, are guaranteed by closed-loop systems whose controllers satisfy such inequalities. Many works in the literature design such controllers as the solution to a state-dependent quadratic program (QP) whose constraints are precisely the inequalities. When the input dimension and number of constraints are high, computing a solution of this QP in real time can become computationally burdensome. Additionally, the solution of such optimization problems is not smooth in general, which can degrade the performance of the system. This paper provides a novel method to design a smooth controller that satisfies an arbitrary number of affine constraints. This why we refer to it as a universal formula for control. The controller is given at every state as the minimizer of a strictly convex function. To avoid computing the minimizer of such function in real time, we introduce a method based on neural networks (NN) to approximate the controller. Remarkably, this NN can be used to solve the controller design problem for any task with less than a fixed input dimension and number of affine constraints, and is completely independent of the state dimension. Additionally, we show that the NN-based controller only needs to be trained with datapoints from a compact set in the state space, which significantly simplifies the training process. Various simulations showcase the performance of the proposed solution, and also show that the NN-based controller can be used to warmstart an optimization scheme that refines the approximation of the true controller in real time, significantly reducing the computational cost compared to a generic initialization.
@@ 1 Introduction
Recent progress in learning-based control has underscored the need to integrate formal control-theoretic guarantees—such as stability, robustness, safety, and planning—into nonlinear dynamical systems, particularly in high-dimensional, uncertain, or data-driven regimes. This need has prompted increasing interest in the intersection of machine learning and control, where rigorous methods can enhance the reliability and safety of autonomous systems, robotics, and reinforcement learning algorithms. Nonlinear control theory provides a unique design framework to ensure that controlled dynamical systems exhibit desired properties of interest. When considering stability, Control Lyapunov Functions (CLFs) (Sontag, 1998, 1983) provide a powerful tool to achieve the stabilization of nonlinear control systems. For control-affine systems, CLFs prescribe an affine inequality in the system input at every state. By design, controllers that satisfy this inequality, such as the pointwise minimum norm controller (Freeman and Kototovic, 1996) or Sontag’s universal formula (Sontag, 1989), ensure stability of the closed-loop system. More recently, Control Barrier Functions (CBFs) (Ames et al., 2019, 2017; Wieland and Allgöwer, 2007) have been introduced to achieve safety requirements for nonlinear control systems. Similarly to CLFs, CBFs also prescribe affine inequalities in the input at every state for control-affine systems, and various safe control designs leveraging this inequality have been proposed (Ames et al., 2019; Cohen et al., 2023; Li et al., 2024). Affine inequalities in the input also arise if control authority is limited (as is always the case in practice), or if stability needs to be guaranteed in fixed time (Garg and Panagou, 2021). There are also a variety of works in the literature that design controllers in scenarios where both stability and safety objectives must be achieved. For example, the CLF-CBF QP in (Ames et al., 2017; Li and Sun, 2023) include both affine constraints in a quadratic program (QP), whose solution defines the desired controller. Alternatively, (Romdlony and Jayawardhana, 2016) unites a CLF and a CBF into a unique function, called control Lyapunov barrier function (CLBF), and uses Sontag’s universal formula to derive a controller. Other control designs include (Mestres and Cortés, 2023), which use penalty methods to impose one of the inequalities as a hard constraint and the other one as a soft constraint, or (Ong and Cortés, 2019), which defines a formula based on the centroids of the two feasible sets defined by the two inequalities.
@@
Often, it is common to face scenarios where the controller must meet multiple objectives, which are encoded by multiple affine inequalities in the input. As an example, a robot might need to visit a set of regions of interest within a prescribed time using fixed-time CLFs, while simultaneously avoiding collisions with various obstacles and other agents in the environment modeled by CBFs, and ensuring input bound constraints are satisfied. All of the aforementioned works are generally limited to up to two affine inequalities. Although all of the inequalities could be encoded in a QP as in (Ames et al., 2017), closed-form expressions for the solutions of QPs only exist for two constraints (Tan and Dimarogonas, 2024), and are difficult to obtain for an arbitrary number. In fact, to the best of our knowledge, a closed-form or computationally cheap controller satisfying an arbitrary number of state-dependent affine constraints is not available in the literature. This raises intertwined theoretical, practical, and computational challenges. Theoretically, the solution of the QP should be computed at every state, which is impossible in practice. This has led to the use sampled-data implementations of such controllers (Singletary et al., 2020; Taylor et al., 2021; Breeden et al., 2022). Even with a sampled-data implementation, from a computational standpoint, solving the QP in real time becomes burdensome and the mismatch between the computed solution and the actual one may degrade system performance.
@@
The problem of designing such a controller is even more challenging if it is required to be smooth. The use of smooth controllers is motivated by theoretical reasons (e.g., existence and uniqueness of solutions and use in backstepping designs (Taylor et al., 2022)) and practical considerations (e.g., avoidance of chattering behavior in digital platforms). Although the existence of an infinitely-times differentiable controller satisfying an arbitrary number of affine state-dependent constraints is guaranteed by (Ong and Cortés, 2019, Proposition 3.1), a constructive method for finding such a controller is lacking. In fact, optimization-based designs such as the QP in (Ames et al., 2017) can fail to be locally Lipschitz (as exemplified by Robinson’s counterexample (Morris et al., 2015)) or even continuous (Alyaseen et al., 2025), and additional constraint qualifications are needed in order to ensure their regularity (Morris et al., 2015; Mestres et al., 2025).
@@
Our work here is also related to a line of research that uses neural networks (NNs) to approximate the solution of model predictive control (MPC) problems (Chen et al., 2022, 2018; Hertneck et al., 2018; Zhang et al., 2019). In the case of linear dynamics and affine input and state constraints, the solution of such MPC problem is known (Borrelli et al., 2017) to be piecewise affine, but computing it explicitly is difficult because the number of regions grows exponentially with the number of constraints. These works show that an approximate solution computed offline with a NN can be used to warmstart the MPC optimization problem online and considerably speed up its computation.
@@
We study the problem of designing a smooth controller that satisfies an arbitrary number of affine constraints at every point in the state space. In our first contribution we introduce a novel controller that generalizes Sontag’s universal formula and satisfies all the affine inequalities. We show that such controller is smooth, and can be computed as the minimizer of a strictly convex function.
@@
Finally, in our third contribution we use the NN approximations in two safe stabilization tasks, both directly and as a warmstart of an optimization scheme to compute the true controller value. We show that the use of the NN significantly reduces the execution time of the controller compared to other standard controllers in the literature such as the minimum-norm CLF-CBF QP controller (Ames et al., 2017), which need to be computed online by solving an optimization problem at every state.
@@
We denote by $\mathbb{Z}_{>0}$ and $\mathbb{R}$ the set of positive integers and real numbers, respectively. For $N\in\mathbb{Z}_{>0}$, we write $[N]=\{1,\ldots,N\}$. Given $l\in\mathbb{Z}_{>0}$ and $\mathcal{S}\subset\mathbb{R}^{n}$, the set of $l$-times continuously differentiable functions in $\mathcal{S}$ is denoted by $\mathcal{C}^{l}(\mathcal{S})$. Vectors are represented by boldface symbols whereas scalars are represented by non-boldface symbols. The zero vector in $\mathbb{R}^{n}$ is denoted by $\mathbf{0}_{n}$, and the identity matrix of dimension $n\times n$, by $\mathbf{I}_{n}$. Given ${\mathbf{x}}\in\mathbb{R}^{n}$, $\lVert{\mathbf{x}}\rVert$ denotes its Euclidean norm. A function $\beta:\mathbb{R}\to\mathbb{R}$ is of class $\mathcal{K}$ if $\beta(0)=0$ and $\beta$ is strictly increasing. If moreover, $\lim\limits_{t\to\infty}\beta(t)=\infty$ and $\lim\limits_{t\to-\infty}\beta(t)=-\infty$, then $\beta$ is of extended class $\mathcal{K}_{\infty}$. A function $V:\mathbb{R}^{n}\to\mathbb{R}$ is positive definite if $V(\mathbf{0}_{n})=\mathbf{0}_{n}$ and $V({\mathbf{x}})>0$ for all ${\mathbf{x}}\neq\mathbf{0}_{n}$. The ball centered at ${\mathbf{p}}\in\mathbb{R}^{n}$ with radius $r>0$ is denoted by $\mathcal{B}_{r}({\mathbf{p}})$. Let $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a locally Lipschitz vector field and consider the dynamical system $\dot{{\mathbf{x}}}=F({\mathbf{x}})$. Local Lipschitzness of $F$ ensures that, for every initial condition ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$, there exists $T>0$ and a unique trajectory $x:[0,T]\to\mathbb{R}^{n}$ such that $x(0)={\mathbf{x}}_{0}$ and $\dot{x}(t)=F(x(t))$. If all solutions exist for all $t\geq 0$, then we say the dynamical system is forward complete. In this case, we let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ denote the flow map, defined by $\Phi_{t}({\mathbf{x}})=x(t)$, where $x(t)$ is the unique solution of the dynamical system starting at $x(0)={\mathbf{x}}$. Given a forward complete dynamical system, a set $\mathcal{K}\subset\mathbb{R}^{n}$ is (positively) forward invariant if ${\mathbf{x}}\in\mathcal{K}$ implies that $\Phi_{t}({\mathbf{x}})\in\mathcal{K}$ for all $t\geq 0$. A point ${\mathbf{p}}$ is Lyapunov stable if, for every open set $U$ containing ${\mathbf{p}}$, there exists an open set $\tilde{U}$ also containing ${\mathbf{p}}$ such that for all ${\mathbf{x}}\in\tilde{U}$, $\Phi_{t}({\mathbf{x}})\in U$ for all $t\geq 0$. An equilibrium point ${\mathbf{p}}$ is asymptotically stable if it is Lyapunov stable and there exists an open set $\bar{U}$ containing ${\mathbf{p}}$ such that $\Phi_{t}({\mathbf{x}})\to{\mathbf{p}}$ as $t\to\infty$ for all ${\mathbf{x}}\in\bar{U}$.
@@ 2 Problem Statement
(Control barrier function for ensuring safety): Let $\mathcal{C}\subset\mathbb{R}^{n}$ be a safe set of interest, and suppose that $\mathcal{C}$ is given by the 00-superlevel set of a continuously differentiable function $h:\mathbb{R}^{n}\to\mathbb{R}$, i.e., $\mathcal{C}=\{{\mathbf{x}}\in\mathbb{R}^{n}:h({\mathbf{x}})\geq 0\}$. This function is a control barrier function (CBF) of $\mathcal{C}$ if, for every ${\mathbf{x}}\in\mathcal{C}$, there exists ${\mathbf{u}}\in\mathbb{R}^{m}$ such that
@@
Formally, let $N\in\mathbb{Z}_{>0}$, and for $i\in[N]$, let $a_{i}:\mathbb{R}^{n}\to\mathbb{R}$, $b_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ be $\mathcal{C}^{l}(\mathbb{R}^{n})$ functions for some $l\in\mathbb{Z}_{\geq 0}$. Let $\mathcal{F}$ be the set of points where the inequalities $\{a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}{\mathbf{u}}<0\}_{i\in[N]}$ are simultaneously strictly feasible, i.e.,
:=$\mathcal{F}:=\{{\mathbf{x}}\in\mathbb{R}^{n}:\exists{\mathbf{u}}\in\mathbb{R}^{m}\ \ \text{s.t.}\ \ a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}{\mathbf{u}}<0,\;\forall i\in[N]\}.$
@@
$a_{1}({\mathbf{x}})$ $=\nabla V({\mathbf{x}})^{\top}f({\mathbf{x}})+W({\mathbf{x}}),\;$ $b_{1}({\mathbf{x}})$ $=g({\mathbf{x}})^{\top}\nabla V({\mathbf{x}}),$
$a_{2}({\mathbf{x}})$ $=-\nabla h({\mathbf{x}})^{\top}f({\mathbf{x}})-\alpha(h({\mathbf{x}})),\;$ $b_{2}({\mathbf{x}})$ $=-g({\mathbf{x}})^{\top}\nabla h({\mathbf{x}}).$
In this case, the property that $\mathcal{F}\neq\emptyset$ is known as strict compatibility of the CLF and CBF pair, cf. (Ong and Cortés, 2019; Mestres and Cortés, 2024). (Ong and Cortés, 2019, Proposition 3.1) is an extension of Artstein’s Theorem (cf. (Artstein, 1983)) showing that if $\mathcal{F}$ is non-empty, there exists a $\mathcal{C}^{\infty}$ controller $k:\mathcal{F}\to\mathbb{R}^{m}$ such that $a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}k({\mathbf{x}})<0$ for all ${\mathbf{x}}\in\mathcal{F}$ and all $i\in[N]$. The proof of this result, however, is not constructive, which is undesirable for the real-time control of the nonlinear system (1). Hence, in this paper we set out to solve the following problem.

### v2

@@ Abstract
We study the problem of designing a controller that satisfies an arbitrary number of affine inequalities at every point in the state space. This is motivated by the fact that a variety of key control objectives, such as stability, safety, and input saturation, are guaranteed by closed-loop systems whose controllers satisfy such inequalities. Many works in the literature design such controllers as the solution to a state-dependent quadratic program (QP) whose constraints are precisely the inequalities. When the input dimension and number of constraints are high, computing a solution of this QP in real time can become computationally burdensome. Additionally, the solution of such optimization problems is not smooth in general, which can degrade the performance of the system. This paper provides a novel method to design a smooth controller that satisfies an arbitrary number of affine constraints. The controller is given at every state as the minimizer of a strictly convex function. To avoid computing the minimizer of such function in real time, we introduce a method based on neural networks (NN) to approximate the controller. Remarkably, this NN can be used to solve the controller design problem for any task with less than a fixed input dimension and number of affine constraints, and is completely independent of the state dimension. This is why we refer to such NN approximation as a NN-based universal formula for control. Additionally, we show that the NN-based controller only needs to be trained with datapoints from a bounded set in the state space, which significantly simplifies the training process. Various simulations showcase the performance of the proposed solution, and also show that the NN-based controller can be used to warmstart an optimization scheme that refines the approximation of the true controller in real time, significantly reducing the computational cost compared to a generic initialization.
keywords:
Universal formula, control Lyapunov function, control barrier function, neural network
@@ 1 Introduction
Modern autonomous systems, ranging from self-driving vehicles to aerospace systems, are usually subject to a variety of operational requirements. These include, but are not limited to, trajectory tracking, disturbance rejection, stabilization, and guaranteeing state and input constraints. Although there exist technical tools to design controllers that satisfy each of these specifications, combining them into a single controller that meets all of them is often challenging. For example, control Lyapunov (respectively, barrier) functions define state-dependent input constraints that guarantee stability (respectively, safety) of the underlying control system. However, designing a smooth control law that satisfies all such constraints and can be computed efficiently in real time is significantly difficult. Addressing this problem is the main motivation for our work here.
1.1 Literature Review
@@
Control Lyapunov Functions (CLFs) [1, 2] provide a powerful tool to achieve the stabilization of nonlinear control systems. For control-affine systems, CLFs prescribe an affine inequality in the system input at every state. By design, controllers that satisfy this inequality, such as the pointwise minimum norm controller [3] or the CLF universal formula [4], ensure stability of the closed-loop system. More recently, Control Barrier Functions (CBFs) [5, 6, 7] have been introduced to achieve safety requirements for nonlinear control systems. Similarly to CLFs, CBFs also prescribe affine inequalities in the input at every state for control-affine systems, and various safe control designs leveraging this inequality have been proposed [5, 8, 9]. Affine inequalities in the input also arise if control authority is limited (as is always the case in practice), or if stability needs to be guaranteed in fixed time [10]. Often, it is common to face scenarios where the controller must meet multiple objectives, which are encoded by multiple affine inequalities in the input. There are also a variety of works in the literature that design controllers in this setting. For example, the CLF-CBF QP in [6, 11] include both affine constraints in a quadratic program (QP), whose solution defines the desired controller. Alternatively, [12] unites a CLF and a CBF into a unique function, called control Lyapunov barrier function (CLBF), and uses the known CLF universal formula to derive a controller. Other control designs include [13], which use penalty methods to impose one of the inequalities as a hard constraint and the other one as a soft constraint, [14], which defines a formula based on the centroids of the two feasible sets defined by the two inequalities, or [15], where different CBF constraints are united into a single one by designing a smooth approximation of the intersection of the different safe sets.
@@
However, the aforementioned approaches only lead to a controller that can be obtained in closed-form or in a computationally tractable manner for a low number of constraints [16]. To the best of our knowledge, there does not exist an approach in the literature for obtaining a controller with such properties for an arbitrary number of state-dependent affine constraints. This problem raises intertwined theoretical, practical, and computational challenges. Theoretically, the solution of the QP should be computed at every state, which is impossible in practice. This has led to the use of sampled-data implementations of such controllers [17, 18, 19]. Even with a sampled-data implementation, from a computational standpoint, solving the QP in real time becomes burdensome and the mismatch between the computed solution and the actual one may degrade system performance.
@@
The problem of designing such a controller is even more challenging if it is required to be smooth. The use of smooth controllers is motivated by theoretical reasons (e.g., existence and uniqueness of solutions and use in backstepping designs [20]) and practical considerations (e.g., avoidance of chattering behavior in digital platforms). Although the existence of an infinitely-times differentiable controller satisfying an arbitrary number of affine state-dependent constraints is guaranteed by [14, Proposition 3.1], a constructive method for finding such a controller is lacking. In fact, optimization-based designs such as the QP in [6] can fail to be locally Lipschitz (as exemplified by Robinson’s counterexample [21]) or even continuous [22], and additional constraint qualifications are needed in order to ensure their regularity [21, 23].
@@
Our work here is also related to a line of research that uses neural networks (NNs) to approximate the solution of model predictive control (MPC) problems [24, 25, 26, 27]. In the case of linear dynamics and affine input and state constraints, the solution of such MPC problem is known [28] to be piecewise affine, but computing it explicitly is difficult because the number of regions grows exponentially with the number of constraints. These works show that an approximate solution computed offline with a NN can be used to warmstart the MPC optimization problem online and considerably speed up its computation.
@@
We study the problem of designing a smooth controller that satisfies an arbitrary number of affine constraints at every point in the state space. In our first contribution we introduce a novel controller that generalizes the CLF universal formula and satisfies all the affine inequalities. We show that such controller is smooth, and can be computed as the minimizer of a strictly convex function.
@@
Finally, in our third contribution we use the NN approximations in two safe stabilization tasks, both directly and as a warmstart of an optimization scheme to compute the true controller value. We show that the use of the NN significantly reduces the execution time of the controller compared to other standard controllers in the literature such as the minimum-norm CLF-CBF QP controller [6], which need to be computed online by solving an optimization problem at every state.
@@
We denote by $\mathbb{Z}_{>0}$ and $\mathbb{R}$ the set of positive integers and real numbers, respectively. For $N\in\mathbb{Z}_{>0}$, we write $N]=\{1,\ldots,N\}$. Given $l\in\mathbb{Z}_{>0}$ and $\mathcal{S}\subset\mathbb{R}^{n}$, the set of $l$-times continuously differentiable functions in $\mathcal{S}$ is denoted by $\mathcal{C}^{l}(\mathcal{S})$. Vectors are represented by boldface symbols whereas scalars are represented by non-boldface symbols. The zero vector in $\mathbb{R}^{n}$ is denoted by $\mathbf{0}_{n}$, and the identity matrix of dimension $n\times n$, by $\mathbf{I}_{n}$. Given ${\mathbf{x}}\in\mathbb{R}^{n}$, $\lVert{\mathbf{x}}\rVert$ denotes its Euclidean norm. A function $\beta:\mathbb{R}\to\mathbb{R}$ is of class $\mathcal{K}$ if $\beta(0)=0$ and $\beta$ is strictly increasing. If moreover, $\lim\limits_{t\to\infty}\beta(t)=\infty$ and $\lim\limits_{t\to-\infty}\beta(t)=$-$\infty$, then $\beta$ is of extended class $\mathcal{K}_{\infty}$. A function $V:\mathbb{R}^{n}\to\mathbb{R}$ is positive definite if $V(\mathbf{0}_{n})=\mathbf{0}_{n}$ and $V({\mathbf{x}})>0$ for all ${\mathbf{x}}\neq\mathbf{0}_{n}$. The ball centered at ${\mathbf{p}}\in\mathbb{R}^{n}$ with radius $r>0$ is denoted by $\mathcal{B}_{r}({\mathbf{p}})$. Let $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a locally Lipschitz vector field and consider the dynamical system $\dot{{\mathbf{x}}}=F({\mathbf{x}})$. Local Lipschitzness of $F$ ensures that, for every initial condition ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$, there exists $T>0$ and a unique trajectory $x:[0,T]\to\mathbb{R}^{n}$ such that $x(0)={\mathbf{x}}_{0}$ and $\dot{x}(t)=F(x(t))$. If all solutions exist for all $t\geq 0$, then we say the dynamical system is forward complete. In this case, we let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ denote the flow map, defined by $\Phi_{t}({\mathbf{x}})=x(t)$, where $x(t)$ is the unique solution of the dynamical system starting at $x(0)={\mathbf{x}}$. Given a forward complete dynamical system, a set $\mathcal{K}\subset\mathbb{R}^{n}$ is (positively) forward invariant if ${\mathbf{x}}\in\mathcal{K}$ implies that $\Phi_{t}({\mathbf{x}})\in\mathcal{K}$ for all $t\geq 0$. A point $\mathbf{p}$} is Lyapunov stable if, for every open set $U$ containing $\mathbf{p}$}, there exists an open set $\tilde{U}$ also containing $\mathbf{p}$} such that for all ${\mathbf{x}}\in\tilde{U}$, $\Phi_{t}({\mathbf{x}})\in U$ for all $t\geq 0$. An equilibrium point $\mathbf{p}$} is asymptotically stable if it is Lyapunov stable and there exists an open set $\bar{U}$ containing $\mathbf{p}$} such that $\Phi_{t}({\mathbf{x}})\to{\mathbf{p}}$ as $t\to\infty$ for all ${\mathbf{x}}\in\bar{U}$.
@@ 2 Problem Statement
(Control barrier function for ensuring safety): Let $\mathcal{C}\subset\mathbb{R}^{n}$ be a safe set of interest, and suppose that $\mathcal{C}$ is given by the 0-superlevel set of a continuously differentiable function $h:\mathbb{R}^{n}\to\mathbb{R}$, i.e., $\mathcal{C}=\{{\mathbf{x}}\in\mathbb{R}^{n}:h({\mathbf{x}})\geq 0\}$. This function is a control barrier function (CBF) of $\mathcal{C}$ if, for every ${\mathbf{x}}\in\mathcal{C}$, there exists ${\mathbf{u}}\in\mathbb{R}^{m}$ such that
@@
Formally, let $N\in\mathbb{Z}_{>0}$, and for $i\in[N]$, let $a_{i}:\mathbb{R}^{n}\to\mathbb{R}$, $b_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be $\mathcal{C}^{l}(\mathbb{R}^{n})$ functions for some $l\in\mathbb{Z}_{\geq 0}$. Let $\mathcal{F}$ be the set of points where the inequalities $\{a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}{\mathbf{u}}<0\}_{i\in[N]}$ are simultaneously strictly feasible, i.e.,
$\mathcal{F}$ $:=\{{\mathbf{x}}\in\mathbb{R}^{n}:\exists{\mathbf{u}}\in\mathbb{R}^{m}\ \ \text{s.t.}$
$\quad\quad\ a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}{\mathbf{u}}<0,\;\forall i\in[N]\}.$
@@
$a_{1}({\mathbf{x}})$ $=\nabla V({\mathbf{x}})^{\top}f({\mathbf{x}})+W({\mathbf{x}}),\;$
$b_{1}({\mathbf{x}})$ $=g({\mathbf{x}})^{\top}\nabla V({\mathbf{x}}),$
$a_{2}({\mathbf{x}})$ $=-\nabla h({\mathbf{x}})^{\top}f({\mathbf{x}})-\alpha(h({\mathbf{x}})),\;$
$b_{2}({\mathbf{x}})$ $=-g({\mathbf{x}})^{\top}\nabla h({\mathbf{x}}).$
In this case, the property that $\mathcal{F}\neq\emptyset$ is known as strict compatibility of the CLF and CBF pair, cf. [14, 29]. [14, Proposition 3.1] is an extension of Artstein’s Theorem (cf. [30]) showing that if $\mathcal{F}$ is non-empty, there exists a $\mathcal{C}^{\infty}$ controller $k:\mathcal{F}\to\mathbb{R}^{m}$ such that $a_{i}({\mathbf{x}})+b_{i}({\mathbf{x}})^{\top}k({\mathbf{x}})<0$ for all $\mathbf{x}$$\in\mathcal{F}$ and all $i\in[N]$. The proof of this result, however, is not constructive, which is undesirable for the real-time control of the nonlinear system (1). Hence, in this paper we set out to solve the following problem.